# Bridge the Modality and Capacity Gaps in Vision-Language Model Selection

### v1

@@ Abstract
Vision Language Models (VLMs) excel in zero-shot image classification by pairing images with textual category names. The expanding variety of Pre-Trained VLMs enhances the likelihood of identifying a suitable VLM for specific tasks. Thus, a promising zero-shot image classification strategy is selecting the most appropriate Pre-Trained VLM from the VLM Zoo, relying solely on the text data of the target dataset without access to the dataset’s images. In this paper, we analyze two inherent challenges in assessing the ability of a VLM in this Language-Only VLM selection: the “Modality Gap”—the disparity in VLM’s embeddings across two different modalities, making text a less reliable substitute for images; and the “Capability Gap”— the discrepancy between the VLM’s overall ranking and its ranking for target dataset, hindering direct prediction of a model’s dataset-specific performance from its general performance. We propose VLM Selection With gAp Bridging (Swab) to mitigate the negative impact of these two gaps. Swab first adopts optimal transport to capture the relevance between open-source datasets and target dataset with a transportation matrix. It then uses this matrix to transfer useful statistics of VLMs from open-source datasets to the target dataset for bridging those two gaps and enhancing the VLM’s capacity estimation for VLM selection. Experiments across various VLMs and image classification datasets validate Swab’s effectiveness.
Machine Learning, ICML
@@ 1 Introduction
Vision-Language Models (VLMs) (Radford et al., 2021; Jia et al., 2021; Singh et al., 2022; Yuan et al., 2021) have demonstrated impressive image-text matching ability. One notable application of VLMs is zero-shot image classification (Radford et al., 2021; Menon & Vondrick, 2023; Ge et al., 2023; Mao et al., 2023), where VLMs are leveraged to generate image classifiers using only class names directly. This zero-shot approach has shown considerable success in image classification, particularly in scenarios with scarce or no training images (Ma et al., 2021; He et al., 2023).
Figure 1: Paradigm of Language-Only VLM Selection (LOVM). The model selection algorithm uses two types of data, including the open-source datasets (which have image and text data) and the text data from the target dataset, to predict the VLM’s absolute or relative performance on a target dataset. It then selects the most appropriate VLM based on the predicted performance.
@@
Despite the success of VLM in image classification, the performance of a VLM may vary substantially according to the datasets and domains (Fang et al., 2022), making it challenging to use a single model to handle all tasks. Fortunately, many open-source VLMs are available (Ilharco et al., 2021), and these VLMs form a vast VLM Zoo. With different architectures, pre-training datasets, or training methods, these VLMs have different strengths. The diverse pre-trained VLMs increase the likelihood of pinpointing at least one VLM that excels in a given target dataset in most cases. [^1] Throughout this paper, the term “VLM” specifically refers to a pre-trained VLM. So one solution for zero-shot image classification is identifying the most suitable VLMs in the zoo for a target dataset without access to the dataset’s images. This VLM selection is termed as “Language-Only VLM Selection” (LOVM) (Zohar et al., 2023), and the paradigm is illustrated in Figure 1.
@@
Two key types of information are available for LOVM. One is the target dataset’s text data, i.e., names of the target classes. The other is the open-source datasets, collected in the form of images with their corresponding class names. Based on these data, the goal is to estimate a VLM’s zero-shot image classification capacity ranking among the VLM zoo on the target dataset.
@@
LOVM encounters two challenges stemming from the inherent heterogeneity in models and datasets. The first challenge is the Modality Gap across different modal features extracted by a VLM. Since the visual and textual features extracted by VLMs tend to cluster into two distinct groups and have gap vectors between them (Liang et al., 2022), using text data as image proxies to rank VLMs is inaccurate. The second challenge is the Capability Gap between the VLM’s overall ranking and its ranking in the specific target dataset. Owing to the VLM’s performance variation across different datasets, the VLM’s average performance on open-source datasets is hard to reflect its performance on a specific target dataset. Thus, selecting a VLM based solely on its general strength proves to be an ineffective strategy.
@@
In this paper, we propose VLM Selection With gAp Bridging (Swab) to address both gaps. The key idea is to reuse the statistics from open-source datasets to estimate the statistics on the target dataset, which mitigates the negative impact of these two gaps. In particular, Swab first uses optimal transport to calculate the transport matrix based on textual similarity between class names of open-source and target datasets. After applying VLMs on open-source datasets to calculate VLMs’ statistics, i.e., the class-specific modality gap vectors and performance rankings of different VLMs, Swab utilizes these statistics to estimate the same type of statistics on the target dataset. After that, Swab uses the estimated gap vectors to align the features of text data with the features of images from the corresponding category, which bridges the modality gap. The estimated VLMs’ ranking also improves the prediction of their rankings on the target task, bridging the capacity gap. The main contributions of our paper are:
@@
  • We analyze two key challenges in LOVM, which are the modality gap across VLM’s different modal features and the capability gap between the VLM’s overall ranking and its ranking on a specific target dataset.
@@ 2.1 Selecting VLMs from a Model Zoo
Zero-Shot Image Classification of VLM. Assume there is a pre-trained VLM $f=(f^{I},f^{T})$ consisting of an image encoder $f^{I}$ and a text encoder $f^{T}$. Given an image classification dataset $\mathcal{T}$ with $k_{\mathcal{T}}$ class names $C_{\mathcal{T}}=\{c_{1}^{\mathcal{T}},\cdots,c_{k_{\mathcal{T}}}^{\mathcal{T}}\}$, we input the class names $C_{\mathcal{T}}$ (probably with templates like “A photo of {class}”) into the VLM’s text encoder $f^{T}$ to get the image classifiers $\{\hat{\bm{t}}_{j}\}_{j=1}^{k_{\mathcal{T}}}$. Then, given a test image ${\bm{x}}_{i}$, we use the image encoder $f^{I}$ to extract its feature $\hat{{\bm{x}}}_{i}$. Finally, we predict the label via the cosine similarity between the image feature $\hat{{\bm{x}}}_{i}$ and image classifiers $\{\hat{\bm{t}}_{j}\}_{j=1}^{k_{\mathcal{T}}}$. The class with the highest cosine similarity to the image is selected as the predicted class $\hat{y}_{i}$. Equation 1 and Equation 2 describe this zero-shot image classification process.
$\hat{{\bm{x}}}_{i}=f^{I}({\bm{x}}_{i}),\ \hat{\bm{t}}_{j}=f^{T}(c_{j}^{\mathcal{T}}).$ (1)
@@
VLM Zoo. In recent years, there have emerged a large number of (pre-trained) VLMs. Assume a collection of $M$ VLMs constitute a VLM Zoo $\mathcal{M}$:
=$\mathcal{M}=\left\{f_{m}=\left(f_{m}^{I},f_{m}^{T}\right)\right\}_{m=1}^{M}.$ (3)
@@
The capability of $f_{m}$ is determined by three key factors: the model architecture (e.g., Transformer (Vaswani et al., 2017), ConvNeXt (Liu et al., 2022)), the pre-trained dataset (e.g., LAION-400M (Schuhmann et al., 2021), MS-COCO (Lin et al., 2014)), and the training method (e.g., contrastive loss (Radford et al., 2021), caption loss (Yu et al., 2022)). Combinations of these factors result in “good and diverse” VLMs in $\mathcal{M}$. Given a dataset $\mathcal{T}$, it is probable to find a suitable VLM from the VLM zoo with high zero-shot image classification performance on $\mathcal{T}$.
Language-Only VLM Selection (LOVM). Rather than using the images from the target dataset, LOVM focuses on the zero-shot scenario where only the text data such as class names $C_{\mathcal{T}}$ from the target dataset are available for VLM selection. Besides, we can obtain some open-source image classification datasets $\mathcal{S}$. The set of class names in $\mathcal{S}$ is $C_{\mathcal{S}}=\{c_{1}^{\mathcal{S}},\cdots,c_{k_{\mathcal{S}}}^{\mathcal{S}}\}$, and the $D_{\mathcal{S}}^{I}$ denote the labelled images in these classes. Given a target task $\mathcal{T}$, the VLM selection method $h$ estimates the zero-shot classification ability of $f_{m}$ based on $C_{\mathcal{T}}$, $C_{\mathcal{S}}$, and $D_{\mathcal{S}}^{I}$ as in Equation 4. Here $m\in[1,\cdots,M]$.
$\hat{r}_{m,\mathcal{T}}=h(f_{m}\;|\;C_{\mathcal{T}},C_{\mathcal{S}},D_{\mathcal{S}}^{I}).$ (4)
$\hat{r}_{m,\mathcal{T}}$ is the predicted ranking of the $m$-th VLM $f_{m}$ on the target dataset $\mathcal{T}$. The higher the ranking, the more probable $f_{m}$ achieves higher zero-shot image classification performance on the target dataset $\mathcal{T}$.
Given the test image set $D_{\mathcal{T}}^{I}$ of the target dataset $\mathcal{T}$ with $|D_{\mathcal{T}}^{I}|$ images, the zero-shot image classification accuracy $p_{m,\mathcal{T}}$ of $f_{m}$ is calculated by:
$p_{m,\mathcal{T}}$ $=\frac{1}{|D_{\mathcal{T}}^{I}|}\underset{(\bm{x}_{i},y_{i})\in D_{\mathcal{T}}^{I}}{\sum}\mathbb{I}\left(y_{i}=f_{m}\left(\bm{x}_{i},C_{\mathcal{T}}\right)\right).$ (5)
@@
$f_{m}\left(\bm{x}_{i},C_{\mathcal{T}}\right)$ represents the predicted class with the same manner as Equation 1 and Equation 2. $\mathbb{I}(\cdot)$ is the indicator function, which outputs 1 if the condition is satisfied, and 0 otherwise. Based on $\{p_{m,\mathcal{T}}\}_{m=1}^{M}$, we obtain the true ranking of $M$ VLMs $\bm{r}_{\mathcal{T}}=[r_{1,\mathcal{T}},\ldots,r_{M,\mathcal{T}}]$ by assigning higher ranking $r$ to models with higher accuracy $p$. Since in practical applications, we can’t obtain the test images set $D_{\mathcal{T}}^{I}$ in advance, the goal of LOVM is to make the predicted ranking $\hat{\bm{r}}=[\hat{r}_{1,\mathcal{T}},\cdots,\hat{r}_{M,\mathcal{T}}]$ be an accurate estimation of the ground truth ranking $\bm{r}_{\mathcal{T}}$ so that the best VLM can be selected from the VLM zoo.
@@
Evaluation of LOVM Methods. We measure the performance of the LOVM algorithm by comparing the ranking similarity between $\bm{r}_{\mathcal{T}}$ and $\hat{\bm{r}}_{\mathcal{T}}$. Specifically, we calculate the Top-5 Recall $R_{5}$ (ranges from 0 to 1) and Kendall’s Rank Correlation $\tau$ (ranges from -1 to 1). The larger these two metrics are, the better the LOVM algorithm is.
@@ 2.2 Possible Paradigms for LOVM
Non-Learning-based LOVM. There are three main paradigms for LOVM. The first paradigm is to neglect the visual encoder and select VLM solely on texts. In detail, we can utilize ChatGPT (Ouyang et al., 2022) to generate auxiliary texts $\tilde{D}_{\mathcal{T}}$ based on class names $C_{\mathcal{T}}$ of $\mathcal{T}$. For example, a generated text for the class “lion” is “A lion is sleeping”. More details are described in the Appendix subsection B.1. These class-specific texts act as substitutes for images, which are referred to as “image proxies” in the following text. Then, whether a VLM $f_{m}$ fits $\mathcal{T}$ could be measured by transferability metrics, e.g., H-Score (Bao et al., 2019) and LogME (You et al., 2021), between the VLM’s text encoder $f_{m}^{T}$ and generated text dataset $\tilde{D}_{\mathcal{T}}$. The second solution relies on the general performance of a certain VLM $f_{m}$. We use some open-source datasets to measure a VLM’s general performance. If $f_{m}$ achieves high zero-shot classification performance over open-source datasets, then it is expected to be competitive on $\mathcal{T}$. The latent assumption is that a VLM’s ranking is relatively consistent across tasks.
@@
Learning-based LOVM. In addition, the ability of a VLM could also be predicted based on a ranker model $f_{R}$. The input of $f_{R}$ is a vector $\bm{s}_{m,\mathcal{T}}$, depicting the dataset-specific representation of $f_{m}$ on $\mathcal{T}$, while the output of $f_{R}$ is the relative/absolute performance $\hat{p}_{m,\mathcal{T}}\in\mathbb{R}$ of $f_{m}$ on $\mathcal{T}$. The $f_{R}$ could be learned on open-source datasets $\mathcal{S}$ (Zhang et al., 2023; Zohar et al., 2023). Due to the availability of both class names $C_{\mathcal{S}}$ and images $D_{\mathcal{S}}^{I}$ in the open-source dataset $\mathcal{S}$ such as ImageNet (Deng et al., 2009), we can calculate each VLM’s representation $\{\bm{s}_{m,n}\}_{m=1,n=1}^{m=M,n=N}$ and true zero-shot image classification accuracy $\{p_{m,n}\}_{m=1,n=1}^{m=M,n=N}$. Here $N$ refers to the number of datasets in $\mathcal{S}$. After constructing the train set, the ranker model $f_{R}$ is learned based on the $\{\bm{s}_{m,n},p_{m,n}\}_{m=1,n=1}^{m=M,n=N}$:
$\min_{f_{R}}\;\sum_{m=1}^{M}\sum_{n=1}^{N}\;\ell(f_{R}(\bm{s}_{m,n}),p_{m,n}).$ (6)
@@
$\ell$ is a loss function that measures the discrepancy between the prediction and the ground-truth. Given $\mathcal{T}$, the learned $f_{R}$ is able to predict the performance $\{\hat{p}_{m,\mathcal{T}}\}_{m=1}^{M}$ over $\{\bm{s}_{m,\mathcal{T}}\}_{m=1}^{M}$ via $\hat{p}_{m,\mathcal{T}}=f_{R}(\bm{s}_{m,\mathcal{T}})$. Finally, we can get the predicted VLMs’ ranking $\hat{\bm{r}}$ based on $\{\hat{p}_{m,\mathcal{T}}\}_{m=1}^{M}$. The representation $\bm{s}_{m,\mathcal{T}}$ is one of the keys in this paradigm, and ModelGPT (Zohar et al., 2023) calculates values $\bm{s}_{m,\mathcal{T}}$ via the capability of a VLM’s text encoder $f_{m}^{T}$.
@@
ModelGPT uses generated text data $\tilde{D}_{\mathcal{T}}$ for $\mathcal{T}$ as substitutes for images to calculate some metrics, which measures the zero-shot ability of $f_{m}$ on unseen images by the classification ability of $f_{m}$ on $\tilde{D}_{\mathcal{T}}$:
$s_{m,\mathcal{T}}^{i}={\rm Metric}_{i}\left(f_{m},\tilde{D}_{\mathcal{T}}\right).$ (7)
@@
Here $\text{Metric}_{i}$ indicates the $i$-th metrics function such as Top-1 Accuracy and F1-Score. For example, the Top-1 Accuracy $s_{m,\mathcal{T}}^{1}$ could be calculated in a similar manner as Equation 2, with the only difference being that the features for the $i$-th text $t_{i}$ in $\tilde{D}_{\mathcal{T}}$ are extracted using a text encoder $f_{m}^{T}$:
$s_{m,\mathcal{T}}^{1}=\frac{1}{|\tilde{D}_{\mathcal{T}}|}\underset{(t_{i},y_{i})\in\tilde{D}_{\mathcal{T}}}{\sum}\mathbb{I}\left(y_{i}=f_{m}(t_{i},C_{\mathcal{T}})\right).$ (8)
@@
Besides, ModelGPT uses some metrics for assessing the features’ quality extracted by the VLM’s text encoder $f_{m}^{T}$. More details are in the Appendix subsection B.2. Moreover, the zero-shot classification performance of $f_{m}$ on open-source datasets $\mathcal{S}$ is also included in $\bm{s}_{m,\mathcal{T}}$ as a general ability measure of $f_{m}$. ModelGPT implements $f_{R}$ as a linear model and sets $\ell$ as the square loss.
Figure 2: Validation Experiments on the Modality Gap and Capacity Gap. (a) Predicted VLMs’ zero-shot image classification accuracy based on auxiliary text data vs. VLM’s true accuracy is based on test images. Each point in the graph represents a model. We calculate Kendall’s tau correlation coefficient $\tau$ and the Mean Absolute Error (MAE) between the predicted accuracy and true accuracy. From the result, we can find that the predicted accuracy calculated by using auxiliary text data poorly aligns with the true accuracy, indicating these text data are ineffective image proxies. (b) We calculate the zero-shot image classification performance rankings of 35 VLMs across 23 datasets. $\bm{r_{i}}$ refers to the vector composed of the $i$-th VLM’s ranking across 23 datasets. We compute the average standard deviations and the mean differences between maximum and minimum values for these VLM’s ranking vectors. The result shows the performance of a VLM varies greatly across different datasets.

### v3

# Bridge the Modality and Capability Gaps in Vision-Language Model Selection

@@ Abstract
Vision Language Models (VLMs) excel in zero-shot image classification by pairing images with textual category names. The expanding variety of Pre-Trained VLMs enhances the likelihood of identifying a suitable VLM for specific tasks. To better reuse the VLM resource and fully leverage its potential on different zero-shot image classification tasks, a promising strategy is selecting appropriate Pre-Trained VLMs from the VLM Zoo, relying solely on the text data of the target dataset without access to the dataset’s images. In this paper, we analyze two inherent challenges in assessing the ability of a VLM in this Language-Only VLM selection: the “Modality Gap”—the disparity in VLM’s embeddings across two different modalities, making text a less reliable substitute for images; and the “Capability Gap”— the discrepancy between the VLM’s overall ranking and its ranking for target dataset, hindering direct prediction of a model’s dataset-specific performance from its general performance. We propose VLM Selection With gAp Bridging (Swab) to mitigate the negative impact of two gaps. Swab first adopts optimal transport to capture the relevance between open-source and target datasets with a transportation matrix. It then uses this matrix to transfer useful statistics of VLMs from open-source datasets to the target dataset for bridging two gaps. By bridging two gaps to obtain better substitutes for test images, Swab can accurately predict the performance ranking of different VLMs on the target task without the need for the dataset’s images. Experiments across various VLMs and image classification datasets validate Swab’s effectiveness.
@@ 1 Introduction
Vision-Language Models (VLMs) [43, 22, 45, 61] have demonstrated impressive image-text matching ability. One notable application of VLMs is zero-shot image classification [43, 37, 14, 35], where VLMs are leveraged to generate image classifiers using only class names directly. This zero-shot approach has shown considerable success in scenarios with scarce or no training images [33, 18].
@@
Despite the success of VLM in image classification, the performance of a VLM may vary substantially according to the datasets and domains [11], making it challenging to use a single model to handle all tasks. Fortunately, many open-source VLMs are available [21], and these VLMs form a vast VLM Zoo. With different architectures, pre-training datasets, or training methods, these VLMs have different strengths. The diverse pre-trained VLMs increase the likelihood of pinpointing at least one VLM that excels in a given target dataset in most cases. [^1] Throughout this paper, the term “VLM” specifically refers to a pre-trained VLM. To more effectively reuse the VLM Zoo across diverse target tasks and unlock its full potential, we need a model selection method to choose suitable VLMs from the VLM Zoo for the target task. However, in scenarios such as zero-shot image classification, many users might not have labeled images for their target tasks, especially those who are not Machine Learning researchers. They prefer to describe their needs in text and use a Model Search Engine to find the most suitable model. So one solution is identifying the most suitable VLMs in the zoo for a target dataset without access to the dataset’s images. This VLM selection is termed as “Language-Only VLM Selection” (LOVM) [64], and the paradigm is illustrated in Figure 1.
@@
Two key types of information are available for LOVM. One is the target dataset’s text data, i.e., names of the target classes and class-related labeled texts generated by LLMs (Details described in Section B.1). The other is the open-source datasets, collected in the form of images with their corresponding class names. Based on these data, the goal is to estimate a VLM’s zero-shot image classification capability ranking among the VLM zoo on the target dataset.
Figure 1: Paradigm of Language-Only VLM Selection (LOVM). Users describe the details of their target tasks in text form, such as class names and image domains. Then, LOVM utilizes this information to generate class-related labeled texts through ChatGPT. These texts serve as substitutes for image samples in subsequent model selection algorithms. The model selection algorithm uses two types of data, including the open-source datasets (which have image and text data) and the text data from the target dataset, to predict the VLM’s absolute or relative performance on a target dataset. It then selects the most appropriate VLM based on the predicted performance.
@@
LOVM encounters two challenges stemming from the inherent heterogeneity in models and datasets. The first challenge is the Modality Gap across different modal features extracted by a VLM. Since the visual and textual features extracted by VLMs tend to cluster into two distinct groups and have gap vectors between them [30], using text data as image proxies to rank VLMs is inaccurate. The second challenge is the Capability Gap between the VLM’s overall ranking and its ranking in the specific dataset. Owing to the VLM’s performance variation across different datasets, the VLM’s average performance on open-source datasets is hard to reflect its performance on a specific target dataset. Thus, selecting a VLM based solely on its general strength may prove to be a less effective strategy.
@@
In this paper, we propose VLM Selection With gAp Bridging (Swab) to address both gaps. The key idea is to reuse VLMs’ statistics from open-source datasets to estimate their statistics on the target dataset, which mitigates the negative impact of these two gaps. In particular, Swab first uses optimal transport to calculate the transport matrix based on textual similarity between class names of open-source and target datasets. After applying VLMs on open-source datasets to calculate VLMs’ statistics, i.e., the class-specific modality gap vectors and performance rankings of different VLMs, Swab utilizes these statistics to estimate the same type of statistics on the target dataset. After that, Swab uses the estimated gap vectors to align the features of texts with the features of images from the corresponding category, which bridges the modality gap. Meanwhile, Swab’s estimated VLMs’ ranking also improves the prediction of their rankings on the target task, bridging the capability gap. The related work is in the appendix C. The main contributions are:
@@
  • We analyze two key challenges in LOVM — the modality gap across VLM’s modal features and the capability gap between the VLM’s overall ranking and its ranking on the target dataset.
@@ 2.1 Selecting VLMs from a Model Zoo
Zero-Shot Image Classification of VLM. Assume there is a pre-trained VLM $f=(f^{I},f^{T})$ consisting of an image encoder $f^{I}$ and a text encoder $f^{T}$. Given an image classification dataset $\mathcal{T}$ with $k_{\mathcal{T}}$ class names $C_{\mathcal{T}}=\{c_{1}^{\mathcal{T}},\cdots,c_{k_{\mathcal{T}}}^{\mathcal{T}}\}$, we input the class names $C_{\mathcal{T}}$ (probably with templates like “A photo of {class}”) into the VLM’s text encoder $f^{T}$ to get the image classifiers $\{\hat{\bm{t}}_{j}\}_{j=1}^{k_{\mathcal{T}}}$. Then, given a test image ${\bm{x}}_{i}$, we use the image encoder $f^{I}$ to extract its feature $\hat{{\bm{x}}}_{i}$. Finally, we predict the label via the cosine similarity between the image feature $\hat{{\bm{x}}}_{i}$ and image classifiers $\{\hat{\bm{t}}_{j}\}_{j=1}^{k_{\mathcal{T}}}$. The class with the highest cosine similarity to the image is selected as the predicted class $\hat{y}_{i}$. Given $\hat{{\bm{x}}}_{i}=f^{I}({\bm{x}}_{i}),\ \hat{\bm{t}}_{j}=f^{T}(c_{j}^{\mathcal{T}})$, Equation 1 describes this zero-shot image classification process:
@@
VLM Zoo. In recent years, a large number of (pre-trained) VLMs have emerged. Assume a collection of $M$ VLMs constitute a VLM Zoo =$\mathcal{M}=\left\{f_{m}=\left(f_{m}^{I},f_{m}^{T}\right)\right\}_{m=1}^{M}$. The capability of $f_{m}$ is determined by three key factors: the model architecture (e.g., Transformer [50], ConvNeXt [32]), the pre-trained dataset (e.g., LAION-400M [44], MS-COCO [31]), and the training method (e.g., contrastive loss [43], caption loss [60]). Combinations of these factors result in “good and diverse” VLMs in $\mathcal{M}$. Given a dataset $\mathcal{T}$, it is probable to find a suitable VLM from the VLM zoo with high zero-shot image classification performance on $\mathcal{T}$.
@@
Language-Only VLM Selection (LOVM). Rather than using images from the target dataset, LOVM focuses on the zero-shot scenario where only the target dataset’s text data, such as its class names $C_{\mathcal{T}}$, are available for VLM selection. Besides, we can obtain some open-source image classification datasets $\mathcal{S}$. The set of class names in $\mathcal{S}$ is $C_{\mathcal{S}}=\{c_{1}^{\mathcal{S}},\cdots,c_{k_{\mathcal{S}}}^{\mathcal{S}}\}$, and the $D_{\mathcal{S}}^{I}$ denote the labelled images in these classes. Given a target task $\mathcal{T}$, the VLM selection method $h$ estimates the zero-shot classification ability of $f_{m}$ based on $C_{\mathcal{T}}$, $C_{\mathcal{S}}$, and $D_{\mathcal{S}}^{I}$ via $\hat{r}_{m}^{\mathcal{T}}=h(f_{m}\;|\;C_{\mathcal{T}},C_{\mathcal{S}},D_{\mathcal{S}}^{I})$, where $m\in[1,\cdots,M]$. $\hat{r}_{m}^{\mathcal{T}}$ is the predicted ranking of the $m$-th VLM $f_{m}$ on $\mathcal{T}$. The higher the ranking, the more probable $f_{m}$ achieves higher zero-shot image classification performance on $\mathcal{T}$. Assuming we can obtain the test image set $D_{\mathcal{T}}^{I}$ of the target dataset $\mathcal{T}$ with $|D_{\mathcal{T}}^{I}|$ images, then we can calculate the zero-shot image classification accuracy $p_{m}^{\mathcal{T}}$ of $f_{m}$ is calculated by $p_{m}^{\mathcal{T}}=\frac{1}{|D_{\mathcal{T}}^{I}|}\underset{(\bm{x}_{i},y_{i})\in D_{\mathcal{T}}^{I}}{\sum}\mathbb{I}\left(y_{i}=f_{m}\left(\bm{x}_{i},C_{\mathcal{T}}\right)\right)$. $f_{m}\left(\bm{x}_{i},C_{\mathcal{T}}\right)$ represents the predicted class with the same manner as Equation 1. $\mathbb{I}(\cdot)$ is the indicator function, which outputs 1 if the condition is satisfied, and 0 otherwise. Based on $\{p_{m}^{\mathcal{T}}\}_{m=1}^{M}$, we obtain the true ranking of $M$ VLMs $\bm{r}^{\mathcal{T}}=[r_{1}^{\mathcal{T}},\ldots,r_{M}^{\mathcal{T}}]$ by assigning higher ranking $r$ to models with higher accuracy $p$. However, in the zero-shot scenario, we can’t obtain the test images set $D_{\mathcal{T}}^{I}$ in advance. Therefore, the goal of LOVM is to make the predicted ranking $\hat{\bm{r}}^{\mathcal{T}}=[\hat{r}_{1}^{\mathcal{T}},\cdots,\hat{r}_{M}^{\mathcal{T}}]$ be an accurate estimation of the ground truth ranking $\bm{r}^{\mathcal{T}}=[r_{1}^{\mathcal{T}},\ldots,r_{M}^{\mathcal{T}}]$ so that the best VLM can be selected.
@@
Evaluation of LOVM Methods. We measure the performance of the LOVM algorithm by comparing the ranking similarity between $\bm{r}^{\mathcal{T}}$ and $\hat{\bm{r}}^{\mathcal{T}}$. Specifically, we calculate the Top-5 Recall $R_{5}$ (ranges from 0 to 1) and Kendall’s Rank Correlation $\tau$ (ranges from -1 to 1). The larger the better.
@@ 2.2 Possible Paradigms for LOVM
Non-Learning-based LOVM. There are three main paradigms for LOVM. The first paradigm is to neglect the visual encoder and select VLM solely on texts. In detail, we can utilize ChatGPT [40] to generate auxiliary texts $\tilde{D}_{\mathcal{T}}$ based on class names $C_{\mathcal{T}}$ of $\mathcal{T}$. More details are described in Section B.1. These class-specific texts act as “image proxies”. Then, whether a VLM $f_{m}$ fits $\mathcal{T}$ could be measured by transferability metrics, e.g., H-Score [2] and LogME [58], between the VLM’s text encoder $f_{m}^{T}$ and generated text dataset $\tilde{D}_{\mathcal{T}}$. The second paradigm relies on the general performance of a certain VLM $f_{m}$. We use open-source datasets to measure a VLM’s general performance. If $f_{m}$ achieves high zero-shot classification performance over open-source datasets, then it is expected to be competitive on $\mathcal{T}$. These methods assume that a VLM’s ranking is relatively consistent across tasks.
@@
Learning-based LOVM. The third paradigm is based on the learning process. In detail, the ability of a VLM could be predicted based on a ranker model $f_{R}$. The input of $f_{R}$ is a vector $\bm{s}_{m}^{\mathcal{T}}$, depicting the dataset-specific representation of $f_{m}$ on $\mathcal{T}$, while the output of $f_{R}$ is the relative/absolute performance $\hat{p}_{m}^{\mathcal{T}}\in\mathbb{R}$ of $f_{m}$ on $\mathcal{T}$. The $f_{R}$ could be learned on open-source datasets $\mathcal{S}$ [63, 64]. Due to the availability of both class names $C_{\mathcal{S}}$ and images $D_{\mathcal{S}}^{I}$ in the open-source dataset $\mathcal{S}$ such as ImageNet [8], we can calculate each VLM’s representation $\{\bm{s}_{m}^{n}\}_{m=1,n=1}^{M,N}$ and true zero-shot image classification accuracy $\{p_{m}^{n}\}_{m=1,n=1}^{M,N}$. Here $N$ refers to the number of datasets in $\mathcal{S}$. After constructing the train set, the ranker model $f_{R}$ is learned based on the $\{\bm{s}_{m}^{n},p_{m}^{n}\}_{m=1,n=1}^{M,N}$:
$\min_{f_{R}}\;\sum_{m=1}^{M}\sum_{n=1}^{N}\;\ell(f_{R}(\bm{s}_{m}^{n}),p_{m}^{n}).$ (2)
@@
$\ell$ is a loss function that measures the discrepancy between the prediction and the ground truth, which can be Mean Squared Error Loss and Huber Loss, among others. Given $\mathcal{T}$, the learned $f_{R}$ is able to predict the performance $\{\hat{p}_{m}^{\mathcal{T}}\}_{m=1}^{M}$ over $\{\bm{s}_{m}^{\mathcal{T}}\}_{m=1}^{M}$ via $\hat{p}_{m}^{\mathcal{T}}=f_{R}(\bm{s}_{m}^{\mathcal{T}})$. Finally, we can get the predicted VLMs’ ranking $\hat{\bm{r}}$ based on $\{\hat{p}_{m}^{\mathcal{T}}\}_{m=1}^{M}$. This approach has similarities with meta-learning [12, 3]. Meta-learning attempts to use data from multiple datasets to learn a model adaptable to the target task, while Learning-based LOVM employs data from multiple datasets to learn a ranker model for selecting the suitable model from a VLM Zoo for the target task. The representation $\bm{s}_{m}^{\mathcal{T}}$ is one of the keys in this paradigm, and ModelGPT [64] calculates values $\bm{s}_{m}^{\mathcal{T}}$ via the capability of a VLM’s text encoder $f_{m}^{T}$.
@@
ModelGPT uses generated text data $\tilde{D}_{\mathcal{T}}$ as substitutes for images to calculate some metrics, which measures the zero-shot ability of $f_{m}$ on unseen images by the classification ability of $f_{m}$ on $\tilde{D}_{\mathcal{T}}$:
$s_{m,i}^{\mathcal{T}}={\rm Metric}_{i}\left(f_{m},\tilde{D}_{\mathcal{T}}\right).$ (3)
@@
Here $\text{Metric}_{i}$ indicates the $i$-th metrics function such as Top-1 Accuracy and F1-Score. For example, the Top-1 Accuracy $s_{m,1}^{\mathcal{T}}$ could be calculated in a similar manner as Equation 1, with the only difference being that the test samples were replaced with text samples $t_{i}$ instead of image samples $\bm{x}_{i}$:
$s_{m,1}^{\mathcal{T}}=\frac{1}{|\tilde{D}_{\mathcal{T}}|}\underset{(t_{i},y_{i})\in\tilde{D}_{\mathcal{T}}}{\sum}\mathbb{I}\left(y_{i}=f_{m}(t_{i},C_{\mathcal{T}})\right).$ (4)
@@
Besides, ModelGPT uses some metrics for assessing the features’ quality extracted by the VLM’s text encoder $f_{m}^{T}$. More details are in the Section B.2. Moreover, the zero-shot classification performance of $f_{m}$ on ImageNet is also included in $\bm{s}_{m}^{\mathcal{T}}$ as a general ability measure of $f_{m}$. ModelGPT implements $f_{R}$ as a simple linear model.